OpenStudy (anonymous):

Anyone up for probability ? how many times must a fair coin be tossed so that the probability of at least one head exceeds 0.9

OpenStudy (anonymous):

btw its bionomial distribution

OpenStudy (anonymous):

Let n be the number of times a coin is tossed. Find an expression for the probability of at least 1 head for n tosses (use complimentary events to do so). That value needs to exceed 0.9. Then find n.

OpenStudy (kc_kennylau):

I would do 1-P(no heads) :p

OpenStudy (anonymous):

darling u just googled this question and found this answer even i saw that answer thousand times on bored of studies

OpenStudy (anonymous):

thnx for trying though lol

OpenStudy (anonymous):

u think @kc_kennylau i tried log too but seems to not work

OpenStudy (anonymous):

1- 1/2 = 1/2 but the answer is 4

OpenStudy (kc_kennylau):

P(at least one head) =1-P(no heads at all) =1-P(0.5^n)

OpenStudy (kc_kennylau):

sorry it should be 1-0.5^n

OpenStudy (anonymous):

yea its quiet tricky prbably gonna ask my teacher on monday then post answer

OpenStudy (kc_kennylau):

\[\large1-0.5^n>0.9\]\[\large0.5^n<0.1\]

OpenStudy (kc_kennylau):

\[\large\log_{0.5}{0.5^n}<\log_{0.5}0.1\]\[\large n<\log_{0.5}0.1\]

OpenStudy (kc_kennylau):

3.32192809488736234787031942948939017586483139302458061205475639581593477660862521585013974335937015509965737171025025182... http://www.wolframalpha.com/input/?t=crmtb01&f=ob&i=log_0.5%200.1

OpenStudy (kc_kennylau):

Round it up and you get \(\Huge4\) :)

OpenStudy (anonymous):

hahah thnx man but to round it off to 4 u need the value at least to be above 3.5

OpenStudy (anonymous):

even i got that answer

OpenStudy (kc_kennylau):

well I said round it \(\large\mbox{up}\) not round it \(\large\mbox{off}\) :)

OpenStudy (kc_kennylau):

sorry brb

OpenStudy (anonymous):

hahah ok i didnt read it properly sorry

OpenStudy (kc_kennylau):

it's not your fault :)

OpenStudy (anonymous):

but thank u so much for ur time

OpenStudy (kc_kennylau):

oh it's my duty :)